OpenStudy (anonymous):

Need help! A conveyor belt moves 450 ft in 3/8 h. 1200?

OpenStudy (anonymous):

(unit rate)

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

3/8 = 0.375 (h) f=450 ft/h 450/0.375=1200

OpenStudy (anonymous):

Thanks! Can you help me with one more?